The construction company rents a bulldozer worth 6800 Eur under a leasing agreement for 2 years with a nominal interest rate of 6.42% per annum. Determine the amount of the monthly payment, if the company pays 2 payments at the beginning and at the end of the lease the tenant offers to buy a bulldozer for 2000 Eur?
1
Expert's answer
2020-12-31T14:42:26-0500

Answer:

6800*2*(6.42\100)=873.12

The nominal interest will be 873.12 Eur.

The total amount to be paid after 2 years will be 6,800+873.12 which is 7,673.12 Eur.

2 years has 24 months.

7,673.12\24=319.71.

The first 2 payments will be 319.71*2 which is 638.38 Eur.

638.38+2000 = 2,638.38.

7,673.12-2,638.38=5,034.74 Eur.

The company has done 2 payments at the beginning. Thus two months will be substracted from 24 to get the remaining months.

24-2= 22.

Thus, the remaining months are 22.

5,034.74\22=228.85 Eur.

Therefore,

The tenant will make monthly payment of 228.85 Eur.